# Familiarity-Aware Evidence Compression for Retrieval-Augmented Generation

### v2

@@ 4.2 Impact of Ensemble Coefficient on Performance and Perplexity
Table 2: Compression rates of compression-based baselines. The highest compression rate is indicated in bold, while the second highest is underlined.
@@
Raw Document …(skip)… The Tales of Hoffmann is a 1951 British Technicolor film adaptation of Jacques Offenbach’s opera “The Tales of Hoffmann", written, produced and directed by the team of Michael Powell and Emeric Pressburger working under the umbrella of their production company, The Archers. Emeric Pressburger 12.429
@@
Raw Document The Chronicle of Philanthropy is a magazine that covers the nonprofit world. …(skip)… It was founded in 1988 by editor Phil Semas and then managing editor Stacy Palmer. …(skip)… Philanthropy (magazine) Philanthropy is a quarterly magazine published by the Philanthropy Roundtable. First published as a newsletter in 1987, "Philanthropy" became a glossy magazine in 1996. Philanthropy 4.856
@@
4.3 Integration of Parametric and Non-parametric Knowledge
As shown in Figure 3, we compare the accuracy of FaviComp with Llama3.2-3B-Instruct and Llama3-8B-Instruct compression-target pairs on $\mathrm{Hits=0}$ and $\mathrm{Hits=1}$ subsets with the top-performing baselines, Zero-shot Summarization and CompAct [^5] We provide results of FaviComp on vaious alpha values in Section B.3. FaviComp outperforms other baselines in the $\mathrm{Hits=0}$ subset while performing comparably with others in the $\mathrm{Hits=1}$ subset. This proves that FaviComp effectively relies on parametric knowledge rather than evidential knowledge when faced with irrelevant evidence, while maintaining similar effectiveness in utilizing evidential knowledge when relevant evidence is present.
Interestingly, even though CompAct generally performs better on the $\mathrm{Hits=1}$ subset compared to Zero-shot Summarization, it underperforms relative to Zero-shot Summarization on the $\mathrm{Hits=0}$ subset. This suggests that the training may have been biased towards utilizing solely evidential knowledge, rather than effectively leveraging both sources in synergy.
4.4 Compression Rate Comparisons
[…]
@@ 5 Case Study
Table 3 presents two examples from HQA to illustrate how FaviComp effectively familiarizes evidence while seamlessly integrating both parametric and non-parametric knowledge during evidence compression. We compare its output with Raw Document, which does not apply any compression, and Zero-shot Summarization.
@@ 6 Related Works
Evidence Compression for RAG. Recent efforts on evidence compression seek to compress retrieved evidence pieces to filter out unnecessary information and retain only the essential context Wang et al. (2023c); Li et al. (2024d); Ke et al. (2024); Xu et al. (2024); Yoon et al. (2024). Most recently, Jiang et al. (2023a) and Cao et al. (2024) conduct token-level or embedding-based compression to preserve only the query-relevant information using a trained compressor. Xu et al. (2024) and Yoon et al. (2024) train a compression model to generate an abstractive summary of the documents by distilling knowledge from larger language models.
While these methods are successful to some extent, they often achieve suboptimal performance because the compressed context may be unfamiliar to the LM used in the downstream task due to differences in the internal knowledge and prompt preferences between the compression and the target model. In contrast, FaviComp proactively compresses the evidence pieces in a way to lower the target model’s perplexity using an ensemble decoding technique without any training, thereby improving the downstream performance.
Parametric and Non-parametric Knowledge in RAG. While there have been studies on the phenomena of LM’s utilization of both parametric and non-parametric knowledge sources, there is a lack of research focused on effectively synergizing both sources. A few of these efforts introduce counterfactual augmentation (Longpre et al., 2021; Fang et al., 2024; Zhang et al., 2024) and causal intervention (Zhou et al., 2023; Wang et al., 2023a) to mitigate knowledge conflict, which, however, requires explicitly knowing the features of the input that causes such conflict. Zhang et al. (2023) seek to address this issue by incorporating LM-generated context into the LM’s input along with the retrieved documents, thereby integrating both sources of knowledge. However, merely concatenating both contexts is a suboptimal solution, as LMs may still show bias toward one source over the other when generating responses (Longpre et al., 2021; Wu et al., 2024). To address this, FaviComp employs ensemble decoding during the evidence compression, ensuring that both types of knowledge are seamlessly fused together to create a consistent context.
@@ 7 Conclusion
In this study, we introduce FaviComp, a training-free, decoding-time evidence compression method designed to enhance RAG performance by consolidating retrieved evidence set to be more familiar to the target model, while seamlessly integrating parametric knowledge. Our extensive experiments validate the effectiveness of FaviComp on open-domain QA tasks, showing significant improvements over recent evidence compression baselines in multiple datasets. Additionally, FaviComp’s model-agnostic nature allows it to be effortlessly incorporated into various RAG workflows without additional training, making it a versatile tool for enhancing LMs in complex tasks.
@@
Although FaviComp exhibits superior performance in RAG compared to the recent evidence compression baselines, it has some limitations. (1) FaviComp consumes approximately twice as much computation compared to methods that only use a compression model since it needs two inferences (compression and target model) during the ensemble decoding. However, it is a training-free strategy that can be easily plugged into any RAG application. (2) Ensemble decoding requires the compression and target model to share the same vocabulary and tokenizer, which can limit the range of compatible models. Nonetheless, recent studies, such as Gu et al. (2024), have introduced techniques to enable model-agnostic ensemble decoding. This implies that there will be a potential direction of incorporating model-agnostic ensemble decoding with our framework to enable more flexible integration of various models, which we leave as future work.
@@ Appendix A Implementation Details
(2) Generated Context: We use the context generation prompt in Table 7 to generate the context.
(3) Zero-shot Summarization: We use the evidence compression prompt in Table 7 to compress the retrieved documents.
@@ B.1 Other Compression and Target Models
We conduct an experiment where we use Llama3 -8B-Instruct and Mistral-7B-Instruct for both compression and target models. The result in Table 5 demonstrates that FaviComp outperforms all other baselines, supplementing the effectiveness shown in Section 4.1.
@@ B.2 Head-to-Head Comparison with RECOMP-abstractive
Since the lower performance of RECOMP-abstractive might possibly be due to the use of smaller base model for compression (T5-large), we conduct a head-to-head experiment on FaviComp and RECOMP-abstractive by using the same base compression model. We construct training data on NQ, TQA, and HQA according to Xu et al. (2024) and finetune Mistral-7B-Instruct on each of the training data. We train for 7 epochs using LoRA with Adam optimizer with a learning rate of 2e-6 and a batch size of 64. We present the evaluation results in Table 4. Even though using larger base model for compression enhances the performance of RECOMP-abstractive to some extent, it still underperforms compared to training-free FaviComp. This underscores that the familiarization during evidence compression and integration of parametric and non-parametric knowledge are more helpful to the downstream generation than relying on a trained model for evidence compression.
@@ B.3 Performance of $\mathrm{Hits=0}$ and $\mathrm{Hits=1}$ on Varying Alpha Values
We evaluate FaviComp’s performance on evidence-relevant ($\mathrm{Hits=1}$) and evidence-irrelevant ($\mathrm{Hits=0}$) subsets by varying $\alpha$ values. Figure 4 shows that $\alpha=0.5$ or $\alpha=0.7$ performs the best on the $\mathrm{Hits=0}$ subset, while performance declines as $\alpha$ deviates further from the value. This pattern in the $\mathrm{Hits=0}$ subset mirrors the overall performance trend, suggesting that appropriately utilizing parametric knowledge when the evidence is irrelevant is crucial to the overall performance. In the $\mathrm{Hits=1}$ subset, performance remains consistent for $\alpha$ values up to 0.5 but decreases significantly when $\alpha$ exceeds 0.5 due to the diminished utilization of the relevant evidential context.
@@ C.1 Evaluation
The evaluation prompt template is shown in Figure 5. For all the evaluations throughout the experiment, we switch the positions of the Question and Context if doing so results in better performance. System prompts and demonstrations used in the evaluations are presented in Table 6 and Table 8, respectively.
@@ C.2 FaviComp
The prompt templates for evidence compression and context generation of FaviComp are presented in Table 7

### v3

@@ 4.4 Compression Rate Comparisons
Table 3: Compression rates of the baselines and FaviComp.
@@
Raw Document The Tales of Hoffmann is a 1951 British Technicolor film adaptation of Jacques Offenbach’s opera “The Tales of Hoffmann", written, produced and directed by the team of Michael Powell and Emeric Pressburger … Emeric Pressburger 12.429
@@
Raw Document The Chronicle of Philanthropy is a magazine that covers the nonprofit world. … It was founded in 1988 by editor Phil … First published as a newsletter in 1987, "Philanthropy" became a glossy magazine in 1996. Philanthropy 4.856
@@ 5 Case Study
Table˜4 presents two examples from HQA to illustrate how FaviComp effectively familiarizes evidence while seamlessly integrating both parametric and non-parametric knowledge during evidence compression. We compare its output with Raw Document, which does not apply any compression, and Zero-shot Summarization.
@@ 6 Related Works
Evidence Compression for RAG. Recent efforts on evidence compression seek to compress retrieved evidence pieces to filter out unnecessary information and retain only the essential context Wang et al. (2023c); Li et al. (2024d); Ke et al. (2024); Xu et al. (2024); Yoon et al. (2024). Most recently, Xu et al. (2024) and Yoon et al. (2024) train a compression model to generate an abstractive summary of the documents by distilling knowledge from larger language models.
While these methods are successful to some extent, they often achieve suboptimal performance because of the discrepancy between the compression model and the target model, leading unfamiliarity of the context. In contrast, FaviComp proactively compresses the evidence pieces in a way to lower the target model’s perplexity using an ensemble decoding technique without any training, thereby improving the downstream performance.
Parametric and Non-parametric Knowledge in RAG. There has been a lack of research focused on effectively combining both sources. A few of these efforts introduce counterfactual augmentation (Longpre et al., 2021; Fang et al., 2024; Zhang et al., 2024) and causal intervention (Zhou et al., 2023; Wang et al., 2023a) to mitigate knowledge conflict, which, however, requires explicitly knowing the features of the input that causes such conflict. Zhang et al. (2023) seek to address this issue by incorporating LM-generated context into the LM’s input along with the retrieved documents, thereby integrating both sources of knowledge. However, merely concatenating both contexts is a suboptimal solution, as LMs may still show bias toward one source over the other when generating responses (Longpre et al., 2021; Wu et al., 2024). To address this, FaviComp employs ensemble decoding during the evidence compression, ensuring that both types of knowledge are seamlessly fused together to create a consistent context.
@@ 7 Conclusion
In this study, we introduce FaviComp, a training-free, inference-time evidence compression method designed to enhance RAG performance by consolidating retrieved evidence set to be more familiar to the target model, while seamlessly integrating parametric knowledge. Our extensive experiments validate the effectiveness of FaviComp on open-domain QA tasks, showing significant improvements over recent evidence compression baselines in multiple datasets. Additionally, FaviComp’s model-agnostic nature allows it to be incorporated into various RAG workflows at inference time, making it a versatile tool for enhancing LMs in complex tasks.
Acknowledgment
We appreciate the reviewers for their insightful comments and suggestions. This work was partly supported by the Amazon Nova Trusted AI Prize, the NSF of the United States Grants ITE 2333736 and OAC 2531126, and the DARPA FoundSci Grant HR00112490370.
@@ Acknowledgment
Although FaviComp exhibits superior performance in RAG compared to the recent evidence compression baselines, it has some limitations. (1) FaviComp consumes approximately twice as much computation compared to methods that only use a compression model since it needs two inferences (compression and target model) during the ensemble decoding. However, it is a training-free strategy that can be easily plugged into any RAG application. We provide insights on the tradeoff between latency and performance in Section˜B.4. (2) Ensemble decoding requires the compression and target model to share the same vocabulary and tokenizer, which can limit the range of compatible models. Nonetheless, recent studies, such as Gu et al. (2024), have introduced techniques to enable model-agnostic ensemble decoding. This implies that there will be a potential direction of incorporating model-agnostic ensemble decoding with our framework to enable more flexible integration of various models, which we leave as future work.
@@ Appendix A Implementation Details
A.1 Generation Configuration
For all the baselines and FaviComp, we use default temperature and top-p values of the compression model during evidence compression and fix the temperature of the target model to 1.0 during evaluation.
A.2 Dataset Statistics
We provide the statistics of the evaluation dataset utilized in our experiments in Table˜6.
A.3 Implementation Details of Baselines
@@
(2) Generated Context: We use the context generation prompt in Table˜10 to generate the context.
(3) Zero-shot Summarization: We use the evidence compression prompt in Table˜10 to compress the retrieved documents.
@@ B.1 Other Compression and Target Models
We conduct an experiment where we use Llama3 -8B-Instruct and Mistral-7B-Instruct for both compression and target models. The result in Table˜8 demonstrates that FaviComp outperforms all other baselines, supplementing the effectiveness shown in Section˜4.1.
@@ B.2 Head-to-Head Comparison with RECOMP-abstractive
Since the lower performance of RECOMP-abstractive might possibly be due to the use of smaller base model for compression (T5-large), we conduct a head-to-head experiment on FaviComp and RECOMP-abstractive by using the same base compression model. We construct training data on NQ, TQA, and HQA according to Xu et al. (2024) and finetune Mistral-7B-Instruct on each of the training data. We train for 7 epochs using LoRA with Adam optimizer with a learning rate of 2e-6 and a batch size of 64. We present the evaluation results in Table˜5. Even though using larger base model for compression enhances the performance of RECOMP-abstractive to some extent, it still underperforms compared to training-free FaviComp. This underscores that the familiarization during evidence compression and integration of parametric and non-parametric knowledge are more helpful to the downstream generation than relying on a trained model for evidence compression.
@@ B.3 Performance of $\mathrm{Hits=0}$ and $\mathrm{Hits=1}$ on Varying Alpha Values
We evaluate FaviComp’s performance on evidence-relevant ($\mathrm{Hits=1}$) and evidence-irrelevant ($\mathrm{Hits=0}$) subsets by varying $\alpha$ values. Figure˜4 shows that $\alpha=0.5$ or $\alpha=0.7$ performs the best on the $\mathrm{Hits=0}$ subset, while performance declines as $\alpha$ deviates further from the value. This pattern in the $\mathrm{Hits=0}$ subset mirrors the overall performance trend, suggesting that appropriately utilizing parametric knowledge when the evidence is irrelevant is crucial to the overall performance. In the $\mathrm{Hits=1}$ subset, performance remains consistent for $\alpha$ values up to 0.5 but decreases significantly when $\alpha$ exceeds 0.5 due to the diminished utilization of the relevant evidential context.
@@
Methods Compression Model Avg latency per sample (s) Performance
Acc F1
RECOMP-abstractive T5-large 0.22 39.3 43.3
CompAct Mistral-7B-Instruct-v0.2 8.72 42.3 46.1
Zero-shot Summarization Llama-3.2-3B-Instruct 3.99 39.4 43.2
FaviComp Llama-3.2-3B-Instruct 6.43 42.8 46.8
Table 7: Latency and of the baselines and FaviComp
B.4 Latency Ablation Study
Table˜7 shows the latency of our method along with other major baselines to provide insights on the trade-offs between accuracy and latency. We used Llama-3-8B-Instruct as the target model and tested on NQ dataset for the experiment. Although there are trade-offs between latency and accuracy across all methods, training-free FaviComp demonstrates lower latency while achieving higher accuracy than CompAct, which is the supervised baseline that previously achieved SOTA performance.
@@ C.1 Evaluation
The evaluation prompt template is shown in Figure˜5. For all the evaluations throughout the experiment, we switch the positions of the Question and Context if doing so results in better performance. System prompts and demonstrations used in the evaluations are presented in Table˜9 and Table˜11, respectively.
@@ C.2 FaviComp
The prompt templates for evidence compression and context generation of FaviComp are presented in Table˜10.